Sir I got 108 marks in jee mains 2017 & my obc rank is 16196 and I m from delhi which branch i can get in dtu and nsit

Grade:12th pass

1 Answers

Vasanth SR
6 years ago
This question is the sum of all the subquestion related to how much marks to get jee main in DTU for CS, MECH, IT, SE, ECE, MCE, BT, EP etc. So I am very happy to answer this question, it's very helpful to DTU aspirants, it's my opinion, I hope after reading my answer you have the same opinion.

I have written many answers about the jee mains marks to get admission in DTU for general, OBC, SC, ST, and PH category. If I wrote specifically for every category(GEN, SC, ST, OBC, PH) and for every branch (CS, IT, SE, MCE, EEE, ECE, EE, BT, EP), therefore answer is very broad So I wrote answer from general category point of view, If you get minimum those marks, you absolutely get a branch in DTU.

The first student admitted to the college of rank was in 2016= 23

He chooses cs branch in DTU.

If you get that rank, you prepare for jee mains extremely hard and also smart work in the study. in jee mains scores more than 300.

The last student admitted in cs branch = 4263

for this rank you must score nearly scores 250 in jee mains.

The last student admitted in college (Last rank in enviornment branch) = 23865

for this you must be score around 180 in jee mains.

after conclusion of this minimum rank in 2016 was 23 and maximum is 23865. If your rank between then, you absoulately get a branch in DTU(for general category).

For OBC students:

minimum rank in this category was 4898 (that student get admission in cs branch).

for want to this rank, you must be score around 220. I am not included board marks beacuse in 2017 board marks are not calculated for merit list of jee mains.

Last rank of OBC students = 153143(he get enviornment branch)

you score 90, you get this rank.

For SC and ST students:

minimum rank =5757(He choose computer science)

maximum rank = 773939 (he choose enviornment science)

So, brillant and innovative minds their in everywhere (not in only general category). Don't judge a student with category and your rank. every student have own potential and special. Just we help him those students to aware your potential and boost them confidence.

don't choose a college brand name, choose by own interset and your criteria.

I hope this helpful for you: I think I cover all points related to marks of jee mains to get admission in dtu, if i am not then suggest me, i edit in answer.